# ChatENV: An Interactive Vision-Language Model for Sensor-Guided Environmental Monitoring and Scenario Simulation

### v1

@@ VI Results & Discussion
Three-Turn Evaluation. Table III shows ChatENV’s performance on a three-turn task that (i) describes Image 1, (ii) describes Image 2, and (iii) explains their differences. Training is conducted using three annotation regimes, ChatGPT, Gemini (each covering 50% of the data), and their union, and evaluation is done over a held-out, fixed test set comprising both sources’ annotations. Additionally, sensor information are passed to the model for each image pair as part of the input prompt. All models share the same Qwen-2.5 backbone and decoder hyperparameters; the only difference lies in the adaptation strategy.
@@
LoRA Vs Linear Probing: Consistently injecting LoRA adapters enhances performance on every measurement relative to the baseline for both the individual and the combined annotation conditions. Importantly, for the combined set, LoRA increases the scores for SBERT and COMET by about +0.30+0.30. Linear probing, however, is competitive and close metrics compared to LoRA, and could potentially surpass it for KCE-F1 scores. Looking through individual sets, LoRA outperforms the probing method and reaches the best score in all five metrics for the ChatGPT-only condition. This indicates that LoRA improves semantic match and contextual understanding, while linear probing can prove to be an effective lightweight option for situations when surface similarity is most important or when resources are scarce.
Two-Turn “What-If” Evaluation. Table III shows What-If performance when the model is fed a single image and sensor data, generates a weather-aware caption, and then responds to a hypothetical question (e.g., “What if more trees are planted?”) by predicting environmental changes like temperature increase or CO2 reduction. Scores are generally consistent with the trend seen in Table III but are typically lower in some metrics, since the model has to imagine the situation as there is no second image, making the task inherently more difficult. In spite of this increased complexity, the KEC-F1 and SBERT scores are still high, which verifies that the model identifies the direction and meaning of important changes accurately, which is the overall aim for the What-If scenario.
@@
Comparison with SOTA Video VLMs. Table IV shows an ablation study comparing ChatENV and two top video models, Video-LLaVA and LLaVA-NeXT-Video architectures specifically tailored for temporal understanding. The two models treat the two input frames as a brief clip and are trained on a three-turn setup, which allows them to memorize directly. This achieves a modest advantage over most aggregated metrics; for example, the best video configuration achieves a COMET of 0.833 compared to 0.763 for ChatENV.
Most importantly, ChatENV still performs and can surpass the video models on certain metrics. It achieves the best BERT-F1 (0.902 vs. 0.818) and KCE-F1 (0.830 vs. 0.758) scores and almost the same ROUGE-L score, which reflects better agreement with human-annotated data despite needing to deduce temporal change under the absence of a sequential video stream.
Base (non-tuned) versions of the video models underperform the ChatENV baseline, emphasizing the role played by domain exposure: Qwen’s pre-training corpus already contains satellite data, while Video-LLaVA and LLaVA-NeXT don’t have aerial data. Upon fine-tuning with our curated dataset, video-model performance increases by at least 75%, demonstrating the way the suggested data pipeline overcomes this domain gap.
From Tables III and IV, it is clear that Gemini annotations consistently beat out ChatGPT annotations on virtually all metrics. Comparing the two models from Table IV further, Video-LLaVA tends to surpass LLaVA-NeXT-Video on ROUGE-L, COMET, and KCE-F1, among others, using fewer samples, for which Video-LLaVA, together with the Gemini annotations, achieves the best overall score.
@@
Table V (LEFT) showcases the effect of having the sensor information in the prompt versus not having them. This experiment was done using the Three-Turn Setting for the ChatENV model, using LoRA finetuning. It is clear that for all the annotation sets, including the sensor data yields higher scores in all metrics, demonstrating the need for sensor data to further explain and describe certain scenes in the images. Additionally, we compared our model against SOTA change detection models, including TEOChat and RS-LLaVA. While TEOChat can process two images simultaneously, it often struggles to describe changes at the scene level, performing better when specific objects are highlighted or when questions focus on natural disasters affecting the area. RS-LLaVA, on the other hand, processes single images similar to ChatENV, but has difficulty linking sensor readings to the corresponding geographical context. In contrast, ChatENV outperforms both models by providing more comprehensive scene-level change descriptions and coherently integrating sensor data into its responses.
TABLE V: Ablation study under the LoRA (three-turn) setting for the ChatENV model. Left: With vs. without sensor data. Right: Comparison across different vision-language models.
@@
Figure 5 provides a qualitative result of the What-If interaction flow. A user first queries the model about an aerial image with its associated sensor readings, then poses a hypothetical (“What if …?”) question. ChatENV’s response (green) closely matches the reference answer derived from the second image (red), where the what-if scenario answer actually happens (using the second image in the dataset). This demonstrates its ability to reason about unseen outcomes. This fidelity underpins practical use-cases such as environmental monitoring and urban-planning simulations, where grounded yet imaginative projections are essential.
@@ VII Conclusion
This paper introduces ChatENV, a novel interactive environmental VLM that jointly infers over satellite images and real-world sensor data. Through the creation of the largest multimodal dataset of its kind, combining pairs of aerial images and rich environmental metadata, with heterogeneous dual-model annotations, ChatENV enables comprehensive and context-aware environmental understanding. The model facilitates fine-grained descriptions for scenes, spatiotemporal comparisons, and “what-if” scenario simulation. Experimental results show ChatENV could output meaningful and humanized descriptions that are very important for urban planning and environmental monitoring.
Future Work: We aim to (i) integrate additional modalities (e.g., SAR, multispectral, socio-economic data), (ii) incorporate memory-augmented agents for long-term reasoning, (iii) support real-time deployment via continual LoRA updates, and (iv) explore joint video–dialogue pretraining. These steps position ChatENV toward becoming a foundation model for predictive environmental intelligence.
@@ Appendix A Annotation Details and Scoring Criteria
These are the full details for scores based on the criteria to manually evaluate the GPT-4 and Gemini 2.0 annotations. These were presented to the annotators (authors) when reviewing the samples. Please note that the total set was split among the annotators, with no overlap due to the subjectivity behind rating the samples. This eliminates the inter-rater overlap.

### v2

@@ VI Results & Discussion
Three-Turn Evaluation. Table III shows ChatENV’s performance on a three-turn task that (i) describes Image 1, (ii) describes Image 2, and (iii) explains the differences between them. Training is conducted using three annotation regimes: ChatGPT-only, Gemini-only (each covering 50% of the data), and their union; evaluation is done over a held-out, fixed test set comprising both sources’ annotations. Additionally, sensor information is passed to the model for each image pair as part of the input prompt. All models share the same Qwen-2.5 backbone and decoder hyper-parameters; the only difference lies in the adaptation strategy.
@@
LoRA Vs. Linear Probing: Consistently injecting LoRA adapters enhances performance on every measurement relative to the baseline for both the individual and combined annotation conditions. Importantly, for the combined set, LoRA increases the scores for SBERT and COMET by about +0.30+0.30. Linear probing, however, is competitive with metrics close to those of LoRA and could potentially surpass it for KCE-F1 scores. Examining the individual sets, LoRA outperforms the probing method and achieves the best score in all five metrics for the ChatGPT-only condition. This indicates that LoRA improves semantic matching and contextual understanding, while linear probing can prove to be an effective lightweight option for situations where surface similarity is most important or when resources are scarce.
Two-Turn “What-If” Evaluation. Table III shows What-If performance when the model is fed a single image and sensor data, generates a weather-aware caption, and then responds to a hypothetical question (e.g., “What if more trees are planted?”) by predicting environmental changes like temperature increase or CO2 reduction. Scores are generally consistent with the trend seen in Table III but are typically lower in some metrics, since the model has to imagine the situation as there is no second image, making the task inherently more difficult. In spite of this increased complexity, the KEC-F1 and SBERT scores are still high, which verifies that the model identifies the direction and meaning of important changes accurately, which is the overall aim for the "What-If" scenario.
@@
Comparison with SOTA Video VLMs. Table IV shows an ablation study comparing ChatENV and two top video models, Video-LLaVA and LLaVA-NeXT-Video architectures specifically tailored for temporal understanding. The two models treat the two input frames as a brief clip and are trained in a three-turn setup, which allows them to memorize directly. This achieves a modest advantage over most aggregated metrics; for example, the best video configuration achieves a COMET of 0.833 compared to 0.763 for ChatENV.
Most importantly, ChatENV still performs well and can surpass the video models on certain metrics. It achieves the best BERT-F1 (0.902 vs. 0.818) and KCE-F1 (0.830 vs. 0.758) scores, and almost the same ROUGE-L score, which reflects better agreement with human-annotated data despite needing to deduce temporal changes in the absence of a sequential video stream.
Base (non-tuned) versions of the video models underperform the ChatENV baseline, emphasizing the role played by domain exposure: Qwen’s pre-training corpus already contains satellite data, while Video-LLaVA and LLaVA-NeXT don’t have aerial data. Upon fine-tuning with our curated dataset, video-model performance increases by at least 75%, demonstrating how the suggested data pipeline overcomes this domain gap.
From Tables III and IV, it is clear that Gemini annotations consistently outperform ChatGPT annotations on virtually all metrics. Comparing the two models from Table IV further, Video-LLaVA tends to surpass LLaVA-NeXT-Video on ROUGE-L, COMET, and KCE-F1, among others, using fewer samples, for which Video-LLaVA, together with the Gemini annotations, achieves the best overall score.
@@
Table V showcases the effect of having the sensor information in the prompt versus not having them. This experiment was done using the Three-Turn Setting for the ChatENV model, using LoRA finetuning. It is clear that for all the annotation sets, including the sensor data yields higher scores in all metrics, demonstrating the need for sensor data to further explain and describe certain scenes in the images. Additionally, we compared our model against SOTA change detection models (under the same setup), including TEOChat and RS-LLaVA seen in Table V. While TEOChat can process two images simultaneously, it often struggles to describe changes at the scene level, performing better when specific objects are highlighted or when questions focus on natural disasters affecting the area. RS-LLaVA, on the other hand, processes single images similarly to ChatENV but has difficulty linking sensor readings to the corresponding geographical context. In contrast, ChatENV outperforms both models by providing more comprehensive scene-level change descriptions and coherently integrating sensor data into its responses.
TABLE V: Ablation study under the LoRA (three-turn) setting for the ChatENV model. Left: With vs. without sensor data. Right: Comparison across different comparative vision-language models.
@@
Figure 5 provides a qualitative result of the What-If interaction flow. A user first queries the model about an aerial image with its associated sensor readings, then poses a hypothetical (“What if …?”) question. ChatENV’s response (green) closely matches the reference answer derived from the second image (red), where the what-if scenario answer actually happens (using the second image in the dataset). This demonstrates its ability to infer unseen outcomes, enabling applications like environmental monitoring and urban-planning simulations where realistic yet imaginative projections are essential.
@@ VII Conclusion
This paper introduces ChatENV, a novel interactive environmental VLM that jointly infers from satellite images and real-world sensor data. Through the creation of the largest multimodal dataset of its kind, which combines pairs of aerial images with rich environmental metadata and heterogeneous dual-model annotations, ChatENV enables comprehensive and context-aware environmental understanding. The model facilitates fine-grained descriptions of scenes, spatiotemporal comparisons, and “what-if” scenario simulations. Experimental results indicate that ChatENV generates informative, human-centered descriptions that are crucial for urban planning and environmental monitoring.
Future Work: We aim to (i) integrate additional modalities (e.g., SAR, multispectral, socio-economic data), (ii) incorporate memory-augmented agents for long-term reasoning, (iii) support real-time deployment via continual LoRA updates, and (iv) explore joint video–dialogue pretraining. These steps position ChatENV toward becoming a foundational model for predictive environmental intelligence.
[Uncaptioned image] Hosam Elgendy is currently a Research Associate in the Machine Learning department at Mohamed bin Zayed University of Artificial Intelligence (MBZUAI) in Abu Dhabi, UAE. He previously received his Master of Science degree in Computer Vision from MBZUAI in 2024. He obtained his Bachelor of Science in Electrical Engineering from the American University of Sharjah (AUS), Sharjah, United Arab Emirates. His research focuses on vision–language models and multi-modal learning systems. He worked extensively on integrating temporal reasoning, conversational AI, and sensor data into remote sensing applications. His interests span computer vision, natural language processing, and multi-modal data fusion, with a focus on scalable model evaluation and deployment.
[Uncaptioned image] Ahmed Sharshar is currently pursuing a PhD in Computer Vision at Mohamed bin Zayed University of Artificial Intelligence (MBZUAI) in Abu Dhabi, UAE. He previously received his Master of Science degree in Computer Vision from MBZUAI. He obtained his Bachelor of Engineering degree in Computer Engineering from the Egypt-Japan University of Science and Technology (E-JUST), Egypt. His research primarily focuses on developing lightweight models and expanding their applications across various domains, such as natural language processing, computer vision, and human-computer interaction. Specifically, he aims to make these models more efficient and accessible, ensuring broader usability and practical deployment.
[Uncaptioned image] Ahmed Hesham is an AI researcher and engineer specializing in Natural Language Processing and speech technologies. He earned his Master’s degree in Natural Language Processing from the Mohamed bin Zayed University of Artificial Intelligence (MBZUAI) in Abu Dhabi, UAE, and his Bachelor’s degree in Computer Engineering from the Egypt-Japan University of Science and Technology (E-JUST), Egypt. His research spans large language models, retrieval-augmented generation, and speech processing, with applications in healthcare, multilingual NLP, and Arabic language technologies. He has published in leading international venues such as INTERSPEECH, IEEE Access, and IJCNN. Ahmed’s work focuses on building efficient, inclusive, and production-ready AI systems, aiming to advance both cutting-edge research and real-world deployment.
[Uncaptioned image] Mohsen Guizani (S’85, M’89, SM’99, F’09) received his B.S. (with distinction) and M.S. degrees in electrical engineering, and M.S. and Ph.D. degrees in computer engineering from Syracuse University, New York, in 1984, 1986, 1987, and 1990, respectively. He is currently a Professor with the Machine Learning Department, Mohamed Bin Zayed University of Artificial Intelligence (MBZUAI), Abu Dhabi, UAE. Previously, he served in different academic and administrative positions at the University of Idaho, Western Michigan University, the University of West Florida, the University of Missouri-Kansas City, the University of Colorado-Boulder, and Syracuse University. His research interests include wireless communications and mobile computing, computer networks, mobile cloud computing, security, and smart grid. He was the Editor-in-Chief of IEEE Network. He serves on the Editorial Boards of several international technical journals, and is the Founder and Editor-in-Chief of the Wireless Communications and Mobile Computing journal (Wiley). He is the author of nine books and more than 500 publications in refereed journals and conferences. He has guest edited a number of Special Issues in IEEE journals and magazines. He has also served as a TPC member, Chair, and General Chair of a number of international conferences.
@@ Appendix A Annotation Details and Scoring Criteria
These are the full details for scores based on the criteria to manually evaluate the GPT-4o and Gemini 2.0 annotations. These were presented to the annotators (authors) when reviewing the samples. Please note that the total set was split among the annotators, with no overlap due to the subjectivity behind rating the samples. This eliminates the inter-rater overlap.